# When do Reforms meet fairness concerns in school admissions?

### v3

@@ Abstract.
We study a series of reforms in school admissions mechanisms motivated, among others, by fairness concerns. Prior to the reforms as well as after, the mechanisms induce blocking students: students who miss schools despite preferring them and having higher priority thereat. We demonstrate that some of these reforms improved fairness by adopting mechanisms with fewer blocking students compared to the preexisting ones, while several reforms did not. In addition, we compared the number of blocking students and the number of students with a profitable misreport.
@@
Acknowledgements: Somouaoga’s work in this project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 890648. The results of this publication reflect only the authors’ view and the commission is not responsible for any use that may be made to the information it contains. Support from the Basic Research Program and the “Priority 2030” Program of the National Research University Higher School of Economics is gratefully acknowledged.
@@ 1. Introduction
Fairness is at the forefront of the concerns that led to the policy changes. The most vivid example is, perhaps, the 2007 major reform in England, which covers 146 local school admissions systems. According to the Secretary of State, Alan Johnson, the aim of the reform was to ‘‘ensure that admission authorities – whether local authorities or schools – operate in a fair way’’ (Department for Education and Skills,, 2007). Among other things, the reform prohibited the practice of giving ‘‘priority to children according to the order of other schools named as preference by their parents,’’ known as the first-preference-first principle. This principle requires that a student who ranks a school higher in her list receives a higher admission priority at this school compared to students who rank it lower. Prior to the reform, as many as one-third of schools in England used this principle.
@@
These examples illustrate unfairness concerns with the old mechanisms: they can induce a matching with a so-called blocking student, that is, a student who missed a school while at least one seat at that school has been assigned to a student with a lower grade or priority or even left empty. The blocking student is entitled to this seat, yet she has not been assigned to it. It is important to note that we define the concept of fairness with respect to true preferences and not reported preferences. A matching with no blocking student is stable and is viewed as a fair outcome as it eliminates “justified envy”, a situation in which a student prefers a school which is assigned to another student with lower admission priority (Abdulkadiroğlu and Sönmez,, 2003). [^1] In general, the relation between stability and fairness is more nuanced, see Romm et al., (2020). Gale and Shapley, (1962) show that for any instance there is a student-optimal stable matching, a matching that every student finds at least as good as any other stable matching. This stable matching can be reached by the student-proposing deferred acceptance algorithm by Gale and Shapley, (1962). We refer to it as Gale-Shapley mechanism.
@@
Apart from the first preference first principle, many mechanisms induce blocking students precisely because they have ranking constraints. In such a mechanism each student is allowed to rank-list only a limited number of schools, typically between 3 and 5 (Pathak and Sönmez,, 2013). Even in New York City, where the ranking constraint is 12 and there are more than 1700 schools, around 25% of students report a complete list of 12 schools, while only 5% report 9, 10, or 11 schools, suggesting that around 20% of students in New York City could not list all acceptable schools (Abdulkadiroğlu et al.,, 2009). Students who missed all their listed schools but could have been admitted to unlisted schools will be dissatisfied with the admissions system and deem it unfair. In this paper, we consider all blocking students, whether it concerns listed schools for which admissions authorities can verify priority violation or unlisted schools which lead to dissatisfaction (see Calsamiglia et al.,, 2010).
In this paper, we pursue two goals. First, we investigate whether the reforms led to more fair matching mechanisms. Second, we investigate the relative importance of the vulnerability to manipulation and fairness concerns in the preexisting mechanisms. We first adopt a fairness criterion introduced by Chen and Kesten, (2017). It is based on the set-inclusion of instances with stable outcomes. One mechanism is more fair by stability than another if it induces a stable matching whenever the second induces a stable matching, and the reverse is not true for at least one instance. Although our investigations show that all reforms but some in England resulted in more fair matching mechanisms by this criterion, we argue that it does not enable a satisfactory explanation. The main reason is that the compared mechanisms might be unstable in a large fraction of instances or real life instances lie in this domain. We indeed illustrate that for many of our compared mechanisms real life instances are likely to be in a domain where the compared mechanisms are unstable. The main contribution of the present paper is a new method of counting the number of blocking students.
A mechanism is more fair by counting (the number of blocking students) than another if for each instance it has weakly fewer number of blocking students than the latter, and fewer in at least one instance. [^2] To our knowledge, this criterion has been first used by Roth and Xing, (1997). Niederle and Roth, (2009), Eriksson and Häggström, (2008) and Doğan and Ehlers, (2021) used the criterion of counting the number of blocking pairs, which does not allow comparisons in our setting (see Remark in section 3).
@@
Our main result for this criterion supports an important kind of reform. Broadly, these reforms involve extending ranking constraints in the Gale-Shapley mechanism. The Gale-Shapley mechanism with an extended constraint is more fair by counting than the restricted counterpart. This took place in Chicago (2010), in Ghana (2007, 2008), in Newcastle (2010), and in Surrey (2010) (Pathak and Sönmez,, 2013). For the remaining reforms, this criterion is too strong. We show that after those reforms the number of blocking students may increase.
This analysis considered fairness separately from strategic issues, which is standard in the literature. But each of the mechanisms that we study might give students incentives to misreport their preferences. This poses a serious methodological difficulty on how to measure the unfairness induced by manipulable mechanisms. On the one hand, when students truthfully report their preferences, the set of blocking students is well understood, but it does not necessarily represent the blocking students of the expected outcomes as some students could have incentives to misreport. On the other hand, when students are well sophisticated, the outcome might not exhibit any blocking pair. To address this difficulty, we develop a setting where some students are sophisticated and best respond while others are truthful (as in Pathak and Sönmez,, 2008). [^3] König et al., (2019) study this setting experimentally by assuming that some students report truthfully and allowing others to misreport. In this setting, unfairness arises from the inability of some students to best respond. We show that some of our results carry over to this new setting. We present this argument in section 4.
@@
We now answer the following question. Was fairness more of a concern compared to vulnerability to manipulation? We focus on blocking students and manipulating students, students who could gain by misreporting their preferences while others are truthful. We show that for any instance the constrained Gale-Shapley mechanism has weakly more blocking students than manipulating students. More precisely, any manipulating student is a blocking student of the mechanism. In contrast, for any instance, the constrained Boston mechanism has weakly more manipulating students than blocking students. More precisely, any blocking student is a manipulating student of the mechanism. For the constrained serial dictatorship mechanism any blocking student is a manipulating student and vice versa. A more subtle relationship between stability and manipulability can be seen in the reform in England. After this reform, the mechanisms in most school districts did not become less manipulable (Bonkoungou and Nesterov,, 2021) and they did not become more fair by stability either (Example 1). However, the reform was successful according to at least one dimension by the following criterion: if the reform disrupted fairness — by producing an unstable matching while it was stable before the reform — the new mechanism is not vulnerable to gaming.
@@
Among strategy-proof and Pareto efficient mechanisms, the Gale’s Top Trading Cycles mechanism (Shapley and Scarf,, 1974) is among the most fair by stability when each school has one seat (Abdulkadiroğlu et al.,, 2020). This result also holds for other fairness comparisons, such as the set of blocking students (Dogan and Ehlers,, 2020) and the set of blocking triplets $(i,j,s)$ – student $i$ blocking the matching of school $s$ with student $j$ (Kwon and Shorrer,, 2019). The result holds for any stability comparison that satisfies few basic properties (Dogan and Ehlers,, 2020).
Among Pareto efficient mechanisms, the Efficiency Adjusted Deferred Acceptance mechanism (EADA) due to Kesten, (2010) is among the most fair in terms of blocking pairs and blocking triplets (Doğan and Ehlers,, 2021; Tang and Zhang,, 2021; Kwon and Shorrer,, 2019). Independent from the present work, Doğan and Ehlers, (2021) also use the fairness by counting criterion to show that among efficient mechanisms EADA is not the most fair by counting, unless the priority profile satisfies few acyclicity conditions.
The first papers that studied constrained mechanisms are Romero-Medina, (1998) and Haeringer and Klijn, (2009). They study the stability of Nash equilibrium outcomes of the game induced by these mechanisms. The most important insight is that the Nash equilibrium outcomes of the constrained Boston mechanism are all stable, while the Nash equilibrium outcomes of the constrained Gale-Shapley may not all be stable. [^4] Ergin and Sönmez, (2006) showed that the Nash equilibrium outcomes of the unconstrained Boston mechanism are stable. Besides, the Nash equilibrium outcomes of the constrained Gale-Shapley are a subset of the Nash equilibrium outcomes of any constrained Gale-Shapley with an extended constraint. Therefore, when the Nash equilibrium outcomes of the constrained Gale-Shapley with an extended constraint are all stable, the Nash equilibrium outcomes of the constrained Gale-Shapley with a restricted constraint are also stable.
The rest of the paper is organized as follows. In Section 2, we introduce the model. In Section 3, we present fairness comparisons and in Section 4, we extend them to strategic settings. In Section 5, we study the relationship between stability and manipulability. We present most of the proofs in the Appendix.
@@ 2. Model
where $P$ and $\succ$ are defined as follows. For each student $i$, $P_{i}$ is a strict preference relation $P_{i}$ over $S\cup\{\emptyset\}$, where $\emptyset$ represents the outside option of being unmatched. For each school $s$, $\succ_{s}$ is a strict priority order over $I$. For each student $i$, let $R_{i}$ denote the ‘‘at least as good as’’ relation associated with $P_{i}$. [^5] That is, for each $s,s^{\prime}\in S\cup\{\emptyset\}$, $s\mathrel{R_{i}}s^{\prime}$ if and only $s\mathrel{P_{i}}s^{\prime}$ or $s=s^{\prime}$. School $s$ is acceptable to student $i$ if $s\mathrel{P_{i}}\emptyset$; and it is unacceptable to student $i$ if $\emptyset\mathrel{P_{i}}s$. We extend the priority order $\succ_{s}$ of each school $s$ over $I$ to the set $2^{I}$ of subsets of students and assume that it is responsive to the priority order over $I$ (Roth,, 1985). By definition, the priority order $\succ_{s}$ over $2^{I}$ is responsive if for any students $i,j\in I$ and any subset $N\subset I\setminus\{i,j\}$ such that $\lvert N\lvert<q_{s}$, (i) $N\cup\{i\}\mathrel{\succ_{s}}N$, and (ii) $N\cup\{i\}\mathrel{\succ_{s}}N\cup\{j\}$ if and only if $i\mathrel{\succ_{s}}j$. Given a proper subset $I^{\prime}\subsetneq I$ of students, we will often write a preference profile as $P=(P_{I^{\prime}},P_{-I^{\prime}})$ to emphasize the components for students in $I^{\prime}$. The tuple $(I,S,P,\succ,q)$ is a school choice problem or simply a problem. We assume that there are more students than schools, that is, $\lvert I\lvert>\lvert S\lvert$. The set of students and the set of schools are fixed throughout the paper, and we denote the (school choice) problem by the triple $(P,\succ,q)$.
@@
A matching $\mu$ is a function $\mu:I\rightarrow S\cup\{\emptyset\}$ such that for each school $s$, $\lvert\mu^{-1}(s)\lvert\leq q_{s}$. We say that student $i$ is matched under $\mu$ if $\mu(i)\neq\emptyset$ and unmatched under $\mu$ if $\mu(i)=\emptyset$. Let $(P,\succ,q)$ be a problem. A matching $\mu$ is individually rational under $P$ if for each student $i$, $\mu(i)\mathrel{R_{i}}\emptyset$. A pair $(i,s)$ of a student and a school blocks the matching $\mu$ under $(P,\succ,q)$ if $s\mathrel{P_{i}}\mu(i)$ and either there is a student $j$ such that $\mu(j)=s$ and $i\mathrel{\succ_{s}}j$ or $\lvert\mu^{-1}(s)\lvert<q_{s}$. Student $i$ is a blocking student for the matching $\mu$ under $(P,\succ,q)$ if there is a school $s$ such that the pair $(i,s)$ blocks $\mu$ under $(P,\succ,q)$. A matching $\mu$ is stable at $(P,\succ,q)$ if it is individually rational under $P$ and has no blocking student. We often drop the problem and refer to a stable matching. A mechanism $\varphi$ is a function which maps each problem to a matching. For each problem $(P,\succ,q)$, let $\varphi_{i}(P,\succ,q)$ denote the component for student $i$. A mechanism $\varphi$ is individually rational if for each problem $(P,\succ,q)$ the matching $\varphi(P,\succ,q)$ is individually rational under $P$. A mechanism $\varphi$ is stable if for each problem $(P,\succ,q)$ the matching $\varphi(P,\succ,q)$ is stable at $(P,\succ,q)$. We often drop the problem and say that a mechanism is stable (at the implicitly assumed problem).
@@ 2.1. Mechanisms
  • Step 1: Each student applies to her most-preferred acceptable school (if any). If a student did not rank any school acceptable, then she remains unmatched. Each school $s$ considers its applicants at the first step denoted as $I^{1}_{s}$ and tentatively accepts $\min(q_{s},\lvert I^{1}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. Let $A^{1}_{s}$ denote the set of students whom school $s$ has tentatively accepted at this step.
  • Step t>1: Each student, who is rejected at step $t-1$, applies to her most-preferred acceptable school among those which have not yet rejected her (if any). If a student does not have any remaining acceptable school, then she remains unmatched. Each school $s$ considers the set $A^{t-1}_{s}\cup I^{t}_{s}$, where $I^{t}_{s}$ are its new applicants at this step, and tentatively accepts $\min(q_{s},\lvert A^{t-1}_{s}\cup I^{t}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. Let $A^{t}_{s}$ denote the set of students whom school $s$ has tentatively accepted at this step.
The algorithm stops when no student is rejected. The tentative acceptances become final at this step. Let $GS$ denote this mechanism.
@@
When schools have the same priority order, we call the Gale-Shapley mechanism the serial dictatorship mechanism. [^6] This is a slight abuse of our definition since the domain of a mechanism is the set of all problems — including problems where schools have different priorities. Let $SD$ denote this mechanism. The outcome of this mechanism can be computed via the following simplified process. Students move in sequence following the common priority order. The first student picks her most preferred acceptable school. The next student picks her most preferred acceptable school among the remaining ones, and so on.
@@
Let $FPF$ denote this mechanism.
@@
  • Step 1: Each student applies to her most-preferred acceptable school (if any). Each school $s$, considers its applicants at the first step denoted as $I^{1}_{s}$ and immediately accepts $\min(q_{s},\lvert I^{1}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. For each school $s$, let $q^{1}_{s}=q_{s}-\min(q_{s},\lvert I^{1}_{s}\lvert)$ denote its remaining capacity after this step.
  • Step t>1: Each student who is rejected at step $t-1$, applies to her most-preferred acceptable school among those which have not yet rejected her (if any). Each school $s$ considers its new applicants $I^{t}_{s}$ at this step and immediately accepts $\min(q^{t-1}_{s},\lvert I^{t}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. For each school $s$, let $q^{t}_{s}=q^{t-1}-\min(q^{t-1}_{s},\lvert I^{t}_{s}\lvert)$ denote its remaining capacity after this step.
The algorithm stops when every student is either accepted at some step or has applied to all of her acceptable schools. Let $\beta$ denote this mechanism.
@@
Constrained mechanisms
Romero-Medina, (1998) and Haeringer and Klijn, (2009) observed that in practice matching mechanisms often have ranking constraints and students are allowed to report only a limited number of schools. This means that schools that are listed below a certain position are not considered. Formally, let $k\in\{1,\ldots,\lvert S\lvert\}$ and let student $i$ have $x$ acceptable schools under $P_{i}$. The truncation after the $k$’th acceptable school (if any) of the preference relation $P_{i}$ is the preference relation $P^{k}_{i}$ with $\min(x,k)$ acceptable schools such that all schools are ordered as in $P_{i}$. Let $P^{k}=(P^{k}_{i})_{i\in I}$. The constrained version $\varphi^{k}$ of the mechanism $\varphi$ is the mechanism that assigns to each problem $(P,\succ,q)$ the matching $\varphi(P^{k},\succ,q)$. That is, $\varphi^{k}(P,\succ,q)=\varphi(P^{k},\succ,q)$.

### v4

@@ Abstract.
We study a series of reforms in school admissions mechanisms motivated, among other reasons, by fairness concerns and vulnerability to manipulation. Before the reforms as well as after, the mechanisms were vulnerable to preference manipulation and induced blocking students: students who miss desired schools despite having higher priority or seats thereat left empty. We demonstrate that some of these reforms improved fairness by adopting mechanisms with fewer blocking students compared to the preexisting ones, while several others did not. We identify preexisting mechanisms where fairness consideration was more of an issue than vulnerability to manipulation and those where it is the reverse.
@@
Acknowledgments: Somouaoga’s work in this project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 890648. The results of this publication reflect only the authors’ view and the commission is not responsible for any use that may be made to the information it contains. Support from the Basic Research Program and the “Priority 2030” Program of the National Research University Higher School of Economics is gratefully acknowledged.
@@ 1. Introduction
Fairness is at the forefront of the concerns that led to the policy changes. The most vivid example is, perhaps, the 2007 major reform in England, which covers 146 local school admissions systems. According to the then-Secretary of State, Alan Johnson, the reform aimed to ‘‘ensure that admission authorities – whether local authorities or schools – operate in a fair way’’ (Department for Education and Skills,, 2007). Among other things, the reform prohibited the practice of giving ‘‘priority to children according to the order of other schools named as preference by their parents,’’ known as the first-preference-first principle. This principle requires that a student who ranks a school higher in her list receives a higher admission priority at this school compared to students who rank it lower. Before the reform, as many as one-third of schools in England used this principle.
@@
These examples illustrate unfairness concerns with the old mechanisms: they can induce a matching with a so-called blocking student, that is, a student who missed a school while at least one seat at that school has been assigned to a student with a lower grade or priority or even left empty. The blocking student is entitled to this seat, yet she has not been assigned to it. It is important to note that we define the concept of fairness concerning true preferences and not reported preferences. A matching with no blocking student is stable and is viewed as a fair outcome as it eliminates “justified envy”, a situation in which a student prefers a school that is assigned to another student with lower admission priority (Abdulkadiroğlu and Sönmez,, 2003). [^1] In general, the relation between stability and fairness is more nuanced, see Romm et al., (2020). Gale and Shapley, (1962) show that for any instance there is a student-optimal stable matching, a matching that every student finds at least as good as any other stable matching. This stable matching can be reached by the student-proposing deferred acceptance algorithm by Gale and Shapley, (1962). We refer to it as Gale-Shapley mechanism.
@@
Apart from the first preference first principle, many mechanisms induce blocking students precisely because they have ranking constraints. In such a mechanism each student is allowed to rank-list only a limited number of schools, typically between 3 and 5 (Pathak and Sönmez,, 2013). Even in New York City, where the ranking constraint is 12 and there are more than 1700 schools, around 25% of students report a complete list of 12 schools, while only 5% report 9, 10, or 11 schools, suggesting that around 20% of students in New York City could not list all acceptable schools (Abdulkadiroğlu et al.,, 2009). Students who missed all their listed schools but could have been admitted to unlisted schools will be dissatisfied with the admissions system and deem it unfair. We consider all blocking students, whether it concerns listed schools for which admissions authorities can verify priority violations or unlisted schools that lead to dissatisfaction (see Calsamiglia et al.,, 2010).
Our first goal is to investigate whether the reforms led to more fair matching mechanisms. Our second goal is to investigate the relative importance of vulnerability to manipulation and fairness concerns in the preexisting mechanisms. Chen and Kesten, (2017) propose to compare mechanisms by set inclusion of problems where they produce stable outcomes. However, this notion cannot further distinguish mechanisms in each instance where they are not stable. A finer and complementary notion is useful, in particular, when the compared mechanisms are not stable or a large fraction of instances or real-life instances lie in this domain. We indeed illustrate that for many of our compared mechanisms, real-life instances are likely to be in a domain where the compared mechanisms are not stable.
@@
To address this problem, we count and compare the number of blocking students across mechanisms. In an instance where two mechanisms are not stable, they can still be contrasted using the number of blocking students. [^2] To our knowledge, this criterion has been first used by Roth and Xing, (1997). Niederle and Roth, (2009), Eriksson and Häggström, (2008) and Doğan and Ehlers, (2021) used the criterion of counting the number of blocking pairs, which does not allow comparisons in our setting (see Remark in section 3). Doğan and Ehlers, (2021) and dogan2022 introduced criteria for counting the number of blocking pairs and blocking students. Our investigation led to a result that supports an important kind of reform. Broadly, these reforms involve extending ranking constraints in the Gale-Shapley mechanism. The Gale-Shapley mechanism with a relaxed constraint has weakly fewer blocking students than the restricted counterpart and there are instances where it has fewer blocking students. This took place in Chicago (2010), in Ghana (2007, 2008), in Newcastle (2010), and in Surrey (2010) (Pathak and Sönmez,, 2013). For the remaining reforms, it is not possible to conclude by comparing the number of blocking students. We show that after those reforms the number of blocking students may increase.
@@
We now answer the following question. Was fairness more of a concern compared to vulnerability to manipulation? We focus on blocking students and manipulating students, students who could gain by misreporting their preferences while others are truthful. We show that for any instance the constrained Gale-Shapley mechanism has weakly more blocking students than manipulating students. More precisely, any manipulating student is a blocking student of the mechanism. In contrast, for any instance, the constrained Boston mechanism has weakly more manipulating students than blocking students. More precisely, any blocking student is a manipulating student of the mechanism. For the constrained serial dictatorship mechanism any blocking student is a manipulating student and vice versa. A more subtle relationship between stability and manipulability can be seen in the reform in England. After this reform, the mechanisms in most school districts did not become less manipulable (Bonkoungou and Nesterov,, 2021) and they did not become fairer by stability either (Example 1). However, the reform was successful according to at least one dimension by the following criterion: if the reform disrupted fairness — by producing an unstable matching while it was stable before the reform — the new mechanism is not vulnerable to manipulation.
@@
Among strategy-proof and Pareto efficient mechanisms, the Gale’s Top Trading Cycles mechanism (Shapley and Scarf,, 1974) is among the most fair by stability when each school has one seat (Abdulkadiroğlu et al.,, 2020). This result also holds for other fairness comparisons, such as the set of blocking students (dogan2022) and the set of blocking triplets $(i,j,s)$ – student $i$ blocking the matching of school $s$ with student $j$ (Kwon and Shorrer,, 2019). The result holds for any stability comparison that satisfies a few basic properties (dogan2022).
Among Pareto efficient mechanisms, the Efficiency Adjusted Deferred Acceptance mechanism (EADA) due to Kesten, (2010) is among the most fair in terms of blocking pairs and blocking triplets (Doğan and Ehlers,, 2021; Tang and Zhang,, 2021; Kwon and Shorrer,, 2019). Independent from the present work, Doğan and Ehlers, (2021) also introduced the fairness comparison by counting to show that among efficient mechanisms, EADA is not the most fair by counting unless the priority profile satisfies a few acyclicity conditions.
The first papers that studied constrained mechanisms are Romero-Medina, (1998) and Haeringer and Klijn, (2009). They study the stability of Nash equilibrium outcomes of the game induced by these mechanisms. The most important insight is that the Nash equilibrium outcomes of the constrained Boston mechanism are all stable, while the Nash equilibrium outcomes of the constrained Gale-Shapley may not all be stable. [^3] Ergin and Sönmez, (2006) showed that the Nash equilibrium outcomes of the unconstrained Boston mechanism are stable. Besides, the Nash equilibrium outcomes of the constrained Gale-Shapley are a subset of the Nash equilibrium outcomes of any constrained Gale-Shapley with a relaxed constraint. Therefore, when the Nash equilibrium outcomes of the constrained Gale-Shapley with a relaxed constraint are all stable, the Nash equilibrium outcomes of the constrained Gale-Shapley with a restricted constraint are also stable.
The rest of the paper is organized as follows. In Section 2, we introduce the model. In Section 3, we present fairness comparisons. In Section 4, we study the relationship between stability and manipulability. We present most of the proofs in the Appendix.
@@ 2. Model
where $P$ and $\succ$ are defined as follows. For each student $i$, $P_{i}$ is a strict preference relation $P_{i}$ over $S\cup\{\emptyset\}$, where $\emptyset$ represents the outside option of being unmatched. For each school $s$, $\succ_{s}$ is a strict priority order over $I$. For each student $i$, let $R_{i}$ denote the ‘‘at least as good as’’ relation associated with $P_{i}$. [^4] That is, for each $s,s^{\prime}\in S\cup\{\emptyset\}$, $s\mathrel{R_{i}}s^{\prime}$ if and only $s\mathrel{P_{i}}s^{\prime}$ or $s=s^{\prime}$. School $s$ is acceptable to student $i$ if $s\mathrel{P_{i}}\emptyset$; and it is unacceptable to student $i$ if $\emptyset\mathrel{P_{i}}s$. We extend the priority order $\succ_{s}$ of each school $s$ over $I$ to the set $2^{I}$ of subsets of students and assume that it is responsive to the priority order over $I$ (Roth,, 1985). By definition, the priority order $\succ_{s}$ over $2^{I}$ is responsive if for any students $i,j\in I$ and any subset $N\subset I\setminus\{i,j\}$ such that $\lvert N\lvert<q_{s}$, (i) $N\cup\{i\}\mathrel{\succ_{s}}N$, and (ii) $N\cup\{i\}\mathrel{\succ_{s}}N\cup\{j\}$ if and only if $i\mathrel{\succ_{s}}j$. Let $k\in\{1,\ldots,\lvert S\lvert\}$, and $P_{i}$ a preference relation where student $i$ has $x$ acceptable schools. The truncation of $P_{i}$ after the $k$’th acceptable school (if any) of the preference relation $P_{i}$ is a preference relation with $\min(x,k)$ acceptable schools such that all schools are ordered as in $P_{i}$. Let $P^{k}_{i}$ denote the truncation of $P_{i}$ after the $k$’th acceptable school. Let $P^{k}=(P^{k}_{i})_{i\in I}$. Given a proper subset $I^{\prime}\subsetneq I$ of students, we will often write a preference profile as $P=(P_{I^{\prime}},P_{-I^{\prime}})$ to emphasize the components for students in $I^{\prime}$. The tuple $(I,S,P,\succ,q)$ is a school choice problem or simply a problem. We assume that there are more students than schools, that is, $\lvert I\lvert>\lvert S\lvert$. The set of students and the set of schools are fixed throughout the paper, and we denote the (school choice) problem by the triple $(P,\succ,q)$.
@@
A matching $\mu$ is a function $\mu:I\rightarrow S\cup\{\emptyset\}$ such that for each school $s$, $\lvert\mu^{-1}(s)\lvert\leq q_{s}$. We say that student $i$ is matched under $\mu$ if $\mu(i)\neq\emptyset$ and unmatched under $\mu$ if $\mu(i)=\emptyset$. Let $(P,\succ,q)$ be a problem. A matching $\mu$ is individually rational under $P$ if for each student $i$, $\mu(i)\mathrel{R_{i}}\emptyset$. A pair $(i,s)$ of a student and a school blocks the matching $\mu$ under $(P,\succ,q)$ if $s\mathrel{P_{i}}\mu(i)$ and either there is a student $j$ such that $\mu(j)=s$ and $i\mathrel{\succ_{s}}j$ or $\lvert\mu^{-1}(s)\lvert<q_{s}$. Student $i$ is a blocking student for the matching $\mu$ under $(P,\succ,q)$ if there is a school $s$ such that the pair $(i,s)$ blocks $\mu$ under $(P,\succ,q)$. A matching $\mu$ is stable at $(P,\succ,q)$ if it is individually rational under $P$ and has no blocking student. We often drop the problem and refer to a stable matching. A mechanism $\varphi$ is a function that maps each problem to a matching. For each problem $(P,\succ,q)$, let $\varphi_{i}(P,\succ,q)$ denote the component for student $i$. A mechanism $\varphi$ is individually rational if for each problem $(P,\succ,q)$ the matching $\varphi(P,\succ,q)$ is individually rational under $P$. A mechanism $\varphi$ is stable if for each problem $(P,\succ,q)$ the matching $\varphi(P,\succ,q)$ is stable at $(P,\succ,q)$. We often drop the problem and say that a mechanism is stable (at the implicitly assumed problem).
@@ 2.1. Mechanisms
  • Step 1: Each student applies to her most preferred acceptable school (if any). If a student did not rank any school acceptable, then she remains unmatched. Each school $s$ considers its applicants at the first step denoted as $I^{1}_{s}$ and tentatively accepts $\min(q_{s},\lvert I^{1}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. Let $A^{1}_{s}$ denote the set of students whom school $s$ has tentatively accepted at this step.
  • Step t>1: Each student, who is rejected at step $t-1$, applies to her most preferred acceptable school among those which have not yet rejected her (if any). If a student does not have any remaining acceptable school, then she remains unmatched. Each school $s$ considers the set $A^{t-1}_{s}\cup I^{t}_{s}$, where $I^{t}_{s}$ are its new applicants at this step, and tentatively accepts $\min(q_{s},\lvert A^{t-1}_{s}\cup I^{t}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. Let $A^{t}_{s}$ denote the set of students whose school $s$ has tentatively accepted at this step.
The algorithm stops when no student is rejected. The tentative acceptances become final at this step. Let $GS$ denote this mechanism. Given $k\in\{1,...,\lvert S\lvert\}$, the constrained version $GS^{k}$ of the Gale-Shapley mechanism $GS$ is the mechanism that assigns to each problem $(P,\succ,q)$ the matching $GS(P^{k},\succ,q)$. That is, $GS^{k}(P,\succ,q)=GS(P^{k},\succ,q)$.
@@
When schools have the same priority order, we call the Gale-Shapley mechanism the serial dictatorship mechanism. [^5] This is a slight abuse of our definition since the domain of a mechanism is the set of all problems — including problems where schools have different priorities. Let $SD$ denote this mechanism. The outcome of this mechanism can be computed via the following simplified process. Students move in sequence following the common priority order. The first student picks her most preferred acceptable school. The next student picks her most preferred acceptable school among the remaining ones, and so on. Given $k\in\{1,...,\lvert S\lvert\}$, the constrained version $SD^{k}$ of the Serial Dictatorship mechanism $SD$ is the mechanism that assigns to each problem $(P,\succ,q)$ the matching $SD(P^{k},\succ,q)$. That is, $SD^{k}(P,\succ,q)=SD(P^{k},\succ,q)$.
@@
Let $FPF$ denote this mechanism. Given $k\in\{1,...,\lvert S\lvert\}$, the constrained version $FPF^{k}$ of the First-Preference-First mechanism $FPF$ is the mechanism that assigns to each problem $(P,\succ,q)$ the matching $FPF(P^{k},\succ,q)$. That is, $FPF^{k}(P,\succ,q)=FPF(P^{k},\succ,q)$.
@@
  • Step 1: Each student applies to her most preferred acceptable school (if any). Each school $s$, considers its applicants at the first step denoted as $I^{1}_{s}$ and immediately accepts $\min(q_{s},\lvert I^{1}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. For each school $s$, let $q^{1}_{s}=q_{s}-\min(q_{s},\lvert I^{1}_{s}\lvert)$ denote its remaining capacity after this step.
  • Step t>1: Each student who is rejected at step $t-1$, applies to her most-preferred acceptable school among those who have not yet rejected her (if any). Each school $s$ considers its new applicants $I^{t}_{s}$ at this step and immediately accepts $\min(q^{t-1}_{s},\lvert I^{t}_{s}\lvert)$ of the $\succ_{s}$-highest priority applicants and rejects the remaining ones. For each school $s$, let $q^{t}_{s}=q^{t-1}-\min(q^{t-1}_{s},\lvert I^{t}_{s}\lvert)$ denote its remaining capacity after this step.
The algorithm stops when every student is either accepted at some step or has applied to all of her acceptable schools. Let $\beta$ denote this mechanism. Given $k\in\{1,...,\lvert S\lvert\}$, the constrained version $\beta^{k}$ of the Boston mechanism $\beta$ is the mechanism that assigns to each problem $(P,\succ,q)$ the matching $\beta(P^{k},\succ,q)$. That is, $\beta^{k}(P,\succ,q)=\beta(P^{k},\succ,q)$.